Circular No. 8493

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005ar Further to IAUC 8492, J. Graham and W. Li report the LOSS discovery

of an apparent supernova on unfiltered KAIT images. The new object is located at α = 11^{h}15^{m}19^{s}.03, δ = −3^{o}46′18′′.5 (equinox 2000.0), which is 4′′.6 east and 5′′.6 north of the center of the apparent host galaxy. Available approximate magnitudes for SN 2005ar: Jan. 20.35 UT, [20.0; Feb. 10.35, [18.5; Mar. 10.37, 17.7; 11.36, 17.6.

M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center for Astrophysics; and T. Matheson, National Optical Astronomy Observatory, report that a spectrum (range 350–740 nm) of SN 2005ar, obtained by J. Foster on Mar. 11.31 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), reveals it to be a type-Ib supernova. The spectrum is very similar to spectra of SN 1998dt (Matheson et al. 2001, A.J. 121, 1648), around one week after maximum brightness, and exhibits He I (rest 587.6 nm) in absorption; however, the other He I lines (rest 667.8, 706.5 nm) are weak. Further spectroscopy is encouraged to check and monitor the evolution of the He I lines.

SUPERNOVA 2005as IN NGC 3450 Graham and Li also report the LOSS discovery of another supernova

on unfiltered KAIT images. SN 2005as is located at α = 10^{h}48^{m}03^{s}.52, δ = −20^{o}50′56′′.5 (equinox 2000.0), which is only 0′′.8 north of the center of NGC 3450. Approximate magnitudes for the new object: 2004 Dec. 17.35 UT, [19.5; 2005 Jan. 14.34, [18.0; 21.35, 17.1; Feb. 2.36, 15.9; Mar. 7.31, 17.2. A. V. Filippenko and R. J. Foley, University of California, Berkeley, report that inspection of CCD spectra (range 320–930 nm), obtained on Mar. 11 with the Keck I 10-m telescope (+ LRIS), shows that this supernova is of type Ia, perhaps six weeks past maximum brightness (but severe contamination by host-galaxy light makes this estimate uncertain).

SUPERNOVA 2005aj IN UGC 2411 Filippenko and Foley add that inspection of CCD spectra, obtained

as above, shows that SN 2005aj (cf. IAUC 8488) is of type Ic. The O I 777.4-nm and Ca II near-infrared triplet absorption troughs are very strong. Prominent Na I D absorption produced by Galactic gas (0.17-nm equivalent width) and host-galaxy gas (0.11-nm equivalent width) are qualitatively consistent with the observed red continuum.

2005 March 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green